OpenStudy (anonymous):

Put the numbers in order, then find the median: -1/2, 0.81, 8.9, 1.5, 1.75, 1.32

OpenStudy (anonymous):

-1/2 = -0.5 -0.5 , 0.81 , 1.32 , 1.5 , 1.75 , 8.9 median : (1.32 + 1.5)/2 = 1.41